OpenStudy (anonymous):

Two ways to solve 10(4 + 1/2)

OpenStudy (texaschic101):

10(4 + 1/2) 10 * 4 + 10 * 1/2 --- this is distributive property 40 + 5 = 45 10 * 4 1/2 = 10 * 9/2 = 90/2 = 45

OpenStudy (anonymous):

Thanks